40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 03:08:57 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec DRFPMI_locked.pdf
    Reply  Thu Oct 15 04:14:14 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Thu Oct 15 10:59:12 2015, Koji, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Sat Oct 17 18:55:07 2015, rana, Update, LSC, DRFPMI Locked for 20 sec 
       Reply  Mon Oct 19 15:23:22 2015, ericq, Update, LSC, Longer DRFPMI lock second_DRFPMI_15Oct2015.png
          Reply  Tue Oct 20 11:24:29 2015, ericq, HowTo, LSC, How to DRFPMI 
Message ID: 11692     Entry time: Thu Oct 15 04:14:14 2015     In reply to: 11691
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Locked for 20 sec 

Fast ALS was still a problem tonight. I don't think high frequency ALS noise saturating the PC drive is the issue; I put two 10k poles before the CM board (shooting for just 2-3kHz bandwidth), and the PC drive levels would be stable and low up until the lockloss, which was always conincident with a step in the AO gain.

After working with that for a few hours, we turned back to our more standard locking attempts. First, we dither aligned the PRMI, and then centered the REFL beam on REFL11. It's hard to say for certain, but we may have been a little close to the edge of the PD. The only other thing that differed from Monday's attempts was using 6dB less AO gain when trying the up the overall gain. 

The script now reliably breaks through to stable high powers, we had a handful of pure-RF locks tonight. The digital DARM gain needs tuning, and the CARM bandwidth still isn't at its final state, but these are very tractable. Off the top of my head, the way forward now includes:

  • Set proper final DARM loop shape
  • Set final CARM loop shape
  • Take full sensing matrix
  • Make 1F handoff
  • Set up the CAL model to produce (at least roughly) calibrated spectra
  • measure noise couplings and other fun stuff

Unrelated: I feel that the PRC angular FF may have deteriorated a bit. I'm leaving the PRC locked on carrier to collect data for wiener filter recalculation. 

ELOG V3.1.3-